A suspension of yeast cells is to be filtered at a constant filtration rate of 50 L/min.
The suspension has a solid content of 70 kg/m3 of suspension and the yeast cells
have a bulk density of 800 kg/m3. Laboratory tests indicate that the specific cake
resistance is 40 m/kg and the viscosity of the filtrate is 2.9x10-3 kg/m.s. The filter has
an area of 0.1 m2 and the medium offers negligible resistance. How long can the
filtration rate be maintained before the pressure drop exceeds 1 N/m2? What volume
of cake and filtrate are collected during this time? Note that ρ0 is expressed in mass
of solids per volume of filtrate and NOT total volume of suspension.
1
Expert's answer
2020-07-22T05:50:59-0400

We are given,

Density"=800kg\/m^3"

Specific cake resistance "=40m\/kg"

Viscosity of the filtrate "=2.9\ *10^{-3}kg/m.s"

Area"=0.1m^2"

Pressure"=1N\/m^2"

"\alpha" "=70kg\/m^3"


Formula used.

"t\ =\dfrac{\mu\alpha\rho_0}{2\Delta P}(\dfrac{V}{A})^2"


"R" c"=\alpha\rho_0(\dfrac{V}{A})"


"40\ *0.1\ =70\ *800\ *V\newline 
4\ =56000V\newline V=7.14\ *10^{-5}"


"t\ =\dfrac{2.9*10^{-3} \ \ *70\ \ *800}{2\ *1}\ (\dfrac{7.14*10^{-5}}{0.1})^2"


"=81.2\ \ *(5.1 *10^{-7})"


"=4.1\ *10^{-5}"